# A Peek into Token Bias: Large Language Models Are Not Yet Genuine Reasoners

### v1

@@ D.2 Hypothesis 2
The full experimental results for Hypothesis 3.3 are shown in Figure 5 and Table LABEL:table:h2.
Figure 5: Full experimental results for Hypothesis 3.3 ($n=500$). The perturbed problems alternate the name classic "Linda" to "Bob" in in-context learning exemplars. We run one-shot with and without chain-of-thought prompts. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}$. We conclude that LLMs possess strong token bias to the name "Linda" frequently appearing in classic literature.
@@ D.3 Hypothesis 3
The full experimental results for Hypothesis 3.4 are shown in Figure 6 and Table LABEL:table:h3.
@@ D.4 Hypothesis 4
The full experimental results for Hypothesis 3.5 are shown in Figure 7 and Table LABEL:table:h4.
@@ D.5 Hypothesis 5
The full experimental results for Hypothesis 3.5 (sets-original vs. sets-framing) are shown in Figure 8 and Table LABEL:table:h5.
@@
The full experimental results for Hypothesis 3.5 (framing gold vs. random) are shown in Figure 9 and Table LABEL:table:h5_framing.
@@ D.6 Hypothesis 6
The full experimental results for Hypothesis 3.6 are shown in Figure 10 and Table LABEL:table:h6.

### v2

@@ D.2 Hypothesis 2
The full experimental results for Hypothesis 3.3 are shown in Figure 7, 8 and Table LABEL:table:h2.
@@
Figure 7: Full experimental results for Hypothesis 3.3 ($n=500$). The perturbed problems alternate the name classic "Linda" to "Bob" in in-context learning exemplars. We run one-shot with and without chain-of-thought prompts. To reject the null, we expect ${\color[rgb]{0.921875,0.6796875,0.63671875}n12}>{\color[rgb]{0.28515625,0.5703125,0.6171875}n21}$. We conclude that LLMs possess a strong token bias toward the name "Linda", which frequently appears in classic literature.
Figure 8: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.3.
@@ D.3 Hypothesis 3
The full experimental results for Hypothesis 3.4 are shown in Figure 9, 10 and Table LABEL:table:h3.
@@
Figure 10: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.4.
@@ D.4 Hypothesis 4
The full experimental results for Hypothesis 3.5 are shown in Figure 11, 12 and Table LABEL:table:h4.
@@
Figure 12: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.5.
@@ D.5 Hypothesis 5
The full experimental results for Hypothesis 3.5 (sets-original vs. sets-framing) are shown in Figure 13, 14 and Table LABEL:table:h5.
@@
Figure 14: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.5.
The full experimental results for Hypothesis 3.5 (framing gold vs. random) are shown in Figure 15, 16 and Table LABEL:table:h5_framing.
@@
Figure 16: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.5.
@@ D.6 Hypothesis 6
The full experimental results for Hypothesis 3.6 are shown in Figure 17, 18 and Table LABEL:table:h6.
@@
Figure 18: Comparison of the accuracy scores between the original and perturbed problems for Hypothesis 3.6.